On September 1, 2013, Susan Chao bought a motorcycle for $24,000. She paid $1,100 down and financed the balance with a five-year loan at an APR of 6.6 percent, compounded monthly. She started the monthly payments exactly one month after the purchase (i.e., October 1, 2013). Two years later, at the end of October 2015, Susan got a new job and decided to pay off the loan.

  

If the bank charges her a 1 percent prepayment penalty based on the loan balance, how much must she pay the bank on November 1, 2015? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

fe date B9 cost of motorbike $24,000 $1,100 downpayment amount financed $2,900 4 monthly installment $449.14 loan amortizatio

For formulas and calculations, refer to the image below -

B9 date A. cost of motorbike 24000 downpayment 1100 amount financed =D2-D3 4 %3D monthly installment =PMT(6.6%/12,5*12,-D4,0,

In case you have any query, kindly ask in comments.